OpenStudy (mendicant_bias):

Dealing with Oxidation states: What is the oxidation state of an individual sulfur atom of BaSO4?

OpenStudy (mendicant_bias):

It's been a while since I've done Chemistry and I basically have zero memory of how to do this. Just need somebody to give me a refresher if they have a decent idea.

OpenStudy (anonymous):

No problem. First, you should always put for oxygen -2 or in case of Hydrogen +1. For groups 1, 2 and 3, their oxidation numbers are the same as their group number. ALWAYS. As soon as you do that, you an equation in which the sum of the oxidation numbers is equal to zero. You will have to find for x, that is for sulfur. +2 +x+ (-2*4)=0. Find x. Did you get it?

OpenStudy (mendicant_bias):

Taking a look at this now, Hoslos. Thank you.

OpenStudy (mendicant_bias):

Alright, so Ba is in Group II, so the oxidation number being identical to the group number is applicable; \[Ba = Ba^{2+}\]

OpenStudy (mendicant_bias):

\[Ba^{2+} + S^{x}+O^{4\times -2}= Ba^{2+}+S^{x}+O^{-8}=A^0\]

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

You got it.

OpenStudy (mendicant_bias):

So the oxidation state of S is -6

OpenStudy (mendicant_bias):

Whoops, 6

OpenStudy (mendicant_bias):

?

OpenStudy (mendicant_bias):

Thanks!

OpenStudy (anonymous):

Pardon me, but are you using an iPad ?

OpenStudy (mendicant_bias):

I'm going to try with another example shortly, about to post it in a new thread. No, I am not. Why do you ask that?

OpenStudy (anonymous):

Wanted to show my working to others the way you showed me now and also be able to send attachments. It would have been very advantageous though.

OpenStudy (mendicant_bias):

Oh! Do you know how to use the equation editor?

OpenStudy (mendicant_bias):

I'm confused about what you're asking, heh.

OpenStudy (anonymous):

See when you showed me your equation, having letters and numbers on top of them ? I wish I could do that using my iPad and send via my iPad. Anyways, you can show me your next question after you try.

OpenStudy (mendicant_bias):

Alright. I think you can do it with your iPad, but only if you go through the trouble of connecting an external physical keyboard via bluetooth or something else, or just hand-type out things into the equation editor using the on-screen keyboard. That's how I used the OpenStudy equation editor while on a tablet when I wasn't at home. I figured out the other example applying the same logic, am now looking at some other problems.

OpenStudy (anonymous):

Hmm, thanks for the tip.